OpenStudy (anonymous):

Given segment VW with endpoints V(-2, -6) and W(x+2, y+3), find the coordinates of the midpoint.

OpenStudy (anonymous):

Midpoint Formula The midpoint is found by calculate these two equations for the x and y values. (x, y) of the midpoint = \[(\frac{x1+x2}{2},\frac{y1+y2}{2})\]

OpenStudy (anonymous):

(x1, y1) = (-2, -6) (x2, y2) = (x+2, y+3)

OpenStudy (anonymous):

can you explain it step by step

OpenStudy (anonymous):

alright we know we have to use the midpoint formula because we are asked to find the midpoint from two points of a line.

OpenStudy (anonymous):

The next thing I wrote was to define what x1, x2, y1, y2 were equal to. In this case x1=-2 y1=-6 x2=(x+2) y2 = (y+3)

OpenStudy (anonymous):

So all we have to do is substitute x1, x2, y1, & y2, into the midpoint formula

OpenStudy (anonymous):

So go ahead and try to do that!

OpenStudy (anonymous):

so i put -2 + x+2/2 and -6+y+3/2

OpenStudy (anonymous):

great

OpenStudy (anonymous):

that is your answer

OpenStudy (anonymous):

MY ANSWERS ARE A.(x/2, (y-3)/2) B.(x/2, y/2) C.(x, y-3) D.(x-2, y-3)

OpenStudy (anonymous):

@juantweaver

OpenStudy (anonymous):

you just have to simplify your answer a little bit

OpenStudy (anonymous):

\[(\frac{-2+x+2}{2}, \frac{-6+y+3}{2})\]

OpenStudy (anonymous):

simplify the numerators!

OpenStudy (anonymous):

numerators which are -6 and what not

OpenStudy (anonymous):

just compute -2+x+2=??? AND -6+y+3=???

OpenStudy (anonymous):

would i subtract x and y

OpenStudy (anonymous):

no because they are in separate numerators. just calculate -2+x+2=?? then calculate =6+y+3

OpenStudy (anonymous):

would i get -4y and 9y

OpenStudy (anonymous):

i see you are having some trouble with the algebra. let's work through this -2+x+2=x since -2+2=0 -6+y+3=y-3 since -6+3=-3

OpenStudy (anonymous):

these are just the numerators, we have to recombine them with the denominators from before

OpenStudy (anonymous):

oh okay well since x=0 and y =-3 whats next because on the answers im getting confused

OpenStudy (anonymous):

Next is the final answer

OpenStudy (anonymous):

\[(\frac{x}{2}, \frac{y-3}{2})\]